[Brahms-l] Shift report 20050608 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jun 08 2005 - 23:59:20 EDT
Shifter leader was rk
Shifters were 


1. Beam injected after 9pm, 84x84 high intensity store.
2. Moved FFS to 30deg. Runs:
15310, 15311 - made mistake and included tr.2 instead of tr.6
15312 - 15315 - 400k tr. 3 and 26k tr.6
3. SMaxL was off, L1458A mainframe was off, - switched them on.
D1 magnet tripped twice when ramped to full field, running at 1/8 field now.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jun 8 23:59:32 2005

This archive was generated by hypermail 2.1.8 : Wed Jun 08 2005 - 23:59:44 EDT